How can you find the absolute difference between the left and right riemann sum on a closed interval? Provide an example with your reasoning.


1
Expert's answer
2022-04-18T00:13:23-0400

Suppose that "f(x)" is a function, which is integrable on the interval "[a,b]". Suppose that we have points "x_0=a,x_1,x_2,...,x_{n-1},x_n=b" and "\Delta=x_{i}-x_{i-1}", "i=1,...,n." Then, the left Riemann sum is: "S_l=\sum_{i=0}^{n-1}f(x_i)(x_{i+1}-x_{i})". The right Riemann sum is: "S_r=\sum_{i=0}^{n-1}f(x_{i+1})(x_{i+1}-x_{i})". The absolute difference is: "|S_r-S_l|=\sum_{i=0}^{n-1}(f(x_{i+1})-f(x_{i}))(x_{i+1}-x_{i})". We point out that for small "\Delta" the difference will approach zero, since both sums will tend to the integral "\int_a^bf(x)dx" . Thus, the more points we have, the smaller will be the difference between sums. As an example, we take . "a=0,b=1". "f(x)=x^2". Consider points: "0,\frac14,\frac24,\frac34,1". We have: "S_l=\frac{1}{4}(0^2+(\frac14)^2+(\frac24)^2+(\frac34)^2)=\frac{7}{32},\,S_r=\frac{1}{4}((\frac14)^2+(\frac24)^2+(\frac34)^2+1)=\frac{15}{32}." The difference is "\frac{8}{32}". The value of integral is: "\int_0^1x^2dx=\frac13". For more points the difference between sums will be smaller. Both sums will approach "\frac13."